All electromagnetic waves travel at a speed of 300000000m/s. What is the frequency of light of wavelenght 0.0000006m?
1
Expert's answer
2011-06-29T09:27:39-0400
The frequency of light can be found by the following expression:
f = c/λ = 3 x 108 [m/s[/ 0.6 x 10-6 [m] = 5 x 1012 Hz.